(1) Motor proteins help move cargo inside cells and these proteins typically exert a force of 5 pN over a single stroke that has a 10 nm step size. How many ATP molecules will a motor protein need to consume per stroke?

(2) What are the 4 key components of a basic bio-device. Describe the role of each in a single sentence.

(3) Why is it important for a cell to ensure that the lifetime of most mRNA molecules is short?
Answer #1

(1) Motor protein needs 1 ATP for approximately 80 Å. With a step size of 10 nm = 100 Å the number of ATP required will be

For 80 Å = 1 ATP

for 1 Å = 1/80 ATP

for 100 Å = (1/80)*100 = 1.25ATP molecules will be consumed per stroke.

(3) RNA molecules are individual transcripts of the cell's DNA. They transfer the genetic information of the DNA and provide a template for the production of proteins that regulate all the cell's processes. The small carriers of information are themselves regulated throughout their lifespan, or rather half-life. After being produced, RNA molecules serve as a template for protein production for a limited time, before they are degraded. Knowing the half-life of RNA is significant for scientific studies on the cell cycle. The whole process of cell division depends on the right amount of proteins being available at the right time. If RNAs are not available in the right concentrations at a given phase of the cell cycle, errors occur.